# Pixels Versus Priors: Controlling Knowledge Priors in Vision-Language Models through Visual Counterfacts

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) perform well on tasks such as visual question answering, but it remains unclear whether their reasoning relies more on memorized world knowledge or on the visual information present in the input image. To investigate this, we introduce Visual CounterFact, a new dataset of visually-realistic counterfactuals that put world knowledge priors (e.g, red strawberry) into direct conflict with visual input (e.g, blue strawberry). Using Visual CounterFact, we show that model predictions initially reflect memorized priors, but shift toward visual evidence in mid-to-late layers. This dynamic reveals a competition between the two modalities, with visual input ultimately overriding priors during evaluation. To control this behavior, we propose Pixels Versus Priors (PvP) steering vectors, a mechanism for controlling model outputs toward either world knowledge or visual input through activation-level interventions. On average, PvP successfully shifts 92.5% of color and 74.6% of size predictions from priors to counterfactuals. Together, these findings offer new tools for interpreting and controlling factual behavior in multimodal models. ***Equal contribution. Order determined by coin flip. Code: https://github.com/rsinghlab/pixels_vs_priors
@@ 1 Introduction
Using Visual CounterFact, we find that MLLMs often ignore world knowledge when shown counterfactual images, favoring perceptual input even when prompted for general facts. We then trace where in the forward pass predictions shift from in-weight knowledge (e.g., strawberries are red) to in-context perception (this strawberry is blue), finding that this transition consistently emerges in mid-to-late layers. During this transition, models frequently flip between the two answers, revealing a competition between in-context pixel and in-weights prior information, with pixels often overriding priors in the model’s output. To control this behavior, we use our Pixels Versus Priors steering to control whether a vision-language model relies on knowledge priors or pixel information. PvP is a novel framework to construct steering vectors for vision-language models that control whether the model responds based on memorized knowledge or in-context visual input. Through this steering, we successfully shift an average of 92.5% of color predictions and 74.6% of size predictions from memorized priors to counterfactual answers.
@@ 5.1 MLLMs are Distracted By Counterfactual Images
Model Task CF + “this” WK + “this”
LLaVA-Next Color 81.74 82.09
Size 82.16 89.91
Qwen2-VL Color 83.65 85.22
Size 94.01 98.18
Janus-Pro Color 82.61 86.78
Size 81.70 87.74
Model Task CF + “most” WK + “most”
LLaVA-Next Color 37.91 84.17
Size 32.16 89.51
Qwen2-VL Color 47.48 87.13
Size 23.15 96.64
Janus-Pro Color 43.30 84.52
Size 15.96 86.32
@@ 5.2 Localizing Visual Perception Shifts through Early Decoding
Size
LLaVA-Next Qwen2-VL Janus-Pro
% Samples w/o Flip 43% 77% 66%
% Samples w/ Flip 57% 23% 34%
Avg. ##\## CF $\to$ WK 1.00 0.33 0.80
Avg. ##\## WK $\to$ CF 1.23 0.87 0.92
Color
% Samples w/o Flip 42% 75% 89%
% Samples w/ Flip 58% 25% 11%
Avg. ##\## CF $\to$ WK 0.79 0.18 0.69
Avg. ##\## WK $\to$ CF 1.24 0.88 0.50
Table 2: Flip statistics for size and color attributes with counterfactual images and prompts designed to elicit world knowledge responses. A “flip” occurs when the initially less probable response later exceeds the alternative.
@@
To understand how this visual override emerges during inference, we apply early decoding to track model predictions across layers. This reveals when the model transitions from relying on memorized priors to integrating counterfactual visual input. Figure 4 shows that in the color task when the model is prompted for the world knowledge answer but given a counterfactual image, the probability of the world-knowledge answer rises in mid to late layers, then flips to the counterfactual answer (orange) in the final layers. This “flipping behavior” is most common when the model is prompted to respond with the world knowledge answer and provided with a counterfactual image (Figure 4 Panel B). This delayed integration of visual input leads to errors when the image contradicts memorized associations, matching the results in Table 1.
In contrast, when using “this” prompt (e.g., “what color is this strawberry?”) for identifying the counterfactual attribute, models are confident in the counterfactual answer by the middle layers and rarely flip to the world-knowledge alternative (Figure 4 panel A and Table 2). This confidence is supported by the high inference accuracies seen in Table 1. Despite their confidence in the counterfactual answer, there is still a slight spike in world knowledge answer probability in mid-to-later layers. This slight spike shows that memorized knowledge does not fully disappear from the model, even when presented with contradicting inputs.
@@ 5.3 Controlling World Knowledge Associations with PvP Steering Vectors
LLaVA Color WK $\to$ CF 89.9 [16-19]
CF $\to$ WK 57.2 [11-18]
Size WK $\to$ CF 66.9 [10-18]
CF $\to$ WK 27.6 [14-23]
QwenVL Color WK $\to$ CF 93.6 [18-21]
CF $\to$ WK 56.8 [17-22]
Size WK $\to$ CF 85.5 [17-24]
CF $\to$ WK 52.4 [11-22]
Janus-Pro Color WK $\to$ CF 94.0 [17-20]
CF $\to$ WK 46.3 [16-19]
Size WK $\to$ CF 71.4 [15-18]
CF $\to$ WK 45.2 [17-21]
@@
Table 3 shows the effectiveness of our steering approach across tasks and models, highlighting both the percentage of successful steering of the model and the key layers at which intervention has the highest impact. We apply PvP steering vectors to the set of inputs that the model originally gets incorrect, meaning without PvP-steering, the model performance on this subset of data is 0%. Remarkably, we achieve at least a 90% success rate in flipping model predictions from world knowledge to counterfactual answers in the color task for all models. This demonstrates that MLLMs are not only steerable but highly responsive to targeted interventions, particularly when guided away from strongly encoded world knowledge priors. The size task is more difficult by nature: it requires detecting two objects and reasoning about their size relationship, making it more dependent on deeper, distributed visual processing. This is reflected in lower flip rates and broader intervention windows.
Across models, we observe that the most effective interventions tend to occur within specific mid-to-late layer ranges, typically requiring sustained influence over multiple layers (Table 3). In general, steering the model from world knowledge to a counterfactual (WK $\rightarrow$ CF) demands less intervention than reversing that shift (CF $\rightarrow$ WK), suggesting that overriding memorized priors is easier than restoring them once suppressed by a counterfactual input. Figure 6 visualizes the top two principal components of the last hidden states from LLaVA-Next, comparing representations before and after PvP steering. Each point corresponds to a hidden state obtained from either the prompt “most” (world knowledge) or “this” (pixel-based), with and without PvP. Steering from world knowledge to counterfactual answers (orange line $\rightarrow$ blue line) shifts the representations along a direction closely aligned with the counterfactual-guided prompt. In contrast, steering from pixel-based prompts to world knowledge prior answers (green $\rightarrow$ purple) induces a shift that is less aligned with the original “most” prompt direction, indicating greater representational reorientation is required when steering from counterfactual answers to world knowledge. These patterns align with the results in Table 3, highlighting how both task complexity and the scope of intervention determine when and where the model’s attention can be effectively redirected.
@@ 5.4 Impact of Prompting and Steering Vectors on Attention Mass
Figure 6: First two principal components of sentence embeddings of LLaVA-Next before and after steering from priors to pixels and from pixels to priors.
@@
Among all models, LLaVA-Next shows the strongest and most consistent shifts in attention when interventions are applied, followed by Qwen2-VL and then Janus-Pro. For the size task, steering vectors must act earlier in the network and across more layers to be effective, reflecting the fact that size requires integrating more visual features than color. In contrast, the color task is more localized and easier to influence with a smaller intervention window. These findings show that PvP steering vectors reshape internal attention mechanisms more effectively than prompt changes alone. They offer precise control over how models allocate attention to visual inputs, especially in tasks like size comparison that require broader spatial reasoning. By intervening directly in the model’s representation space, PvP steering enables deeper interpretability and control over MLLM behavior.
@@ Appendix A Creating Visual CounterFact
Next, we describe the creation of Visual CounterFact, a dataset designed to examine how MLLMs use visual input and world knowledge when presented with controlled counterfactual examples. Visual CounterFact is constructed through a four-step pipeline: (1) sourcing objects with known color priors, (2) sampling real-world images that reflect these priors, (3) constructing counterfactual relations that invert these priors, and (4) editing images to produce perceptually plausible counterfactuals. In the color task, we apply pixel-level hue shifts to create new color-object pairs. In the size task, we compose new images that alter the relative scale between objects. These edits are designed to conflict with the model’s linguistically acquired priors while preserving visual realism. Figure 8 illustrates the full pipeline for both perceptual tasks.
@@
(2) Size: Given two objects with the world knowledge relation “[object 1] is larger than [object 2]” we create the counterfactual image reflecting “[object 2] is larger than [object 1]” by combining the segmentation masks of object1 and object2. For the world knowledge image, we resize the masks so that object1 appears significantly larger than object2, and in the counterfactual image, we resize the masks so that object2 appears significantly larger than object1. Specifically, 250$\times$250 pixels versus 80$\times$80 pixels. This size difference visually reflects the intended relation.
@@ Appendix C Attention Mass Details
Color WK $\rightarrow$ CF $21.8\%$ $12.8\%$
Color CF $\rightarrow$ WK $-25.2\%$ $-12.4\%$
@@
Color WK $\rightarrow$ CF $19.5\%$ $11.7\%$
Color CF $\rightarrow$ WK $-11.0\%$ $-10.2\%$
@@ Appendix D Early Decoding
Appendix E PCA Results
Figure 9 shows the principal component analysis (PCA) results for QwenVL and Janus-Pro, complementing the LLaVA-Next results presented in Figure 6. For each model, we project the final hidden state of the last text token into the top two PCA dimensions and visualize the effect of PvP steering vectors. These findings support the asymmetry observed across models and tasks in Table 3. They show that the model’s internal representation space is more easily reoriented away from memorized priors than toward them, further emphasizing the dominance of perception in the presence of conflicting visual input.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) perform well on tasks such as visual question answering, but it remains unclear whether their reasoning relies more on memorized world knowledge or on the visual information present in the input image. To investigate this, we introduce Visual CounterFact, a new dataset of visually-realistic counterfactuals that put world knowledge priors (e.g, red strawberry) into direct conflict with visual input (e.g, blue strawberry). Using Visual CounterFact, we show that model predictions initially reflect memorized priors, but shift toward visual evidence in mid-to-late layers. This dynamic reveals a competition between the two modalities, with visual input ultimately overriding priors during evaluation. To control this behavior, we propose Pixels Versus Priors (PvP) steering vectors, a mechanism for controlling model outputs toward either world knowledge or visual input through activation-level interventions. On average, PvP successfully shifts 99.3% of color and 80.8% of size predictions from priors to counterfactuals. Together, these findings offer new tools for interpreting and controlling factual behavior in multimodal models. ***Equal contribution. Order determined by coin flip.Code: https://github.com/rsinghlab/pixels_vs_priors
@@ 1 Introduction
Using Visual CounterFact, we find that MLLMs often ignore world knowledge when shown counterfactual images, favoring perceptual input even when prompted for general facts. We then trace where in the forward pass predictions shift from in-weight knowledge (e.g., strawberries are red) to in-context perception (this strawberry is blue), finding that this transition consistently emerges in mid-to-late layers. During this transition, models frequently flip between the two answers, revealing a competition between in-context pixel and in-weights prior information, with pixels often overriding priors in the model’s output. To control this behavior, we use our Pixels Versus Priors steering to control whether a vision-language model relies on knowledge priors or pixel information. PvP is a novel framework to construct steering vectors for vision-language models that control whether the model responds based on memorized knowledge or in-context visual input. Through this steering, we successfully shift an average of 99.3% of color predictions and 80.8% of size predictions from memorized priors to counterfactual answers.
@@ 5.1 MLLMs are Distracted By Counterfactual Images
Model Task CF + “this” WK + “this” CF + “most” WK + “most”
LLaVA-Next Color 85.19 87.22 47.26 92.09
Size 82.12 96.42 40.30 95.60
Qwen2-VL Color 84.79 85.40 60.65 90.87
Size 91.20 98.21 28.34 96.29
Janus-Pro Color 86.00 88.03 59.23 90.47
Size 85.14 96.84 18.02 96.01
@@ 5.2 Localizing Visual Perception Shifts through Early Decoding
Size Color
LLaVA-Next Qwen2-VL Janus-Pro LLaVA-Next Qwen2-VL Janus-Pro
% Samples w/o Flip 45% 69% 70% 35% 71% 88%
% Samples w/ Flip 55% 31% 30% 65% 29% 12%
Avg. #\# CF $\to$ WK 1.02 0.47 0.76 0.84 0.18 0.56
Avg. #\# WK $\to$ CF 1.12 0.70 0.86 1.31 0.90 0.56
Table 2: Flip statistics for size and color attributes with counterfactual images and prompts designed to elicit world knowledge responses. A “flip” occurs when the initially less probable response later exceeds the alternative by at least 5%.
@@
To understand how this visual override emerges during inference, we apply early decoding to track model predictions across layers. This reveals when the model transitions from relying on memorized priors to integrating counterfactual visual input. Figure 4 shows that in the color task when the model is prompted for the world knowledge answer but given a counterfactual image, the probability of the world‐knowledge answer rises in mid to late layers, then flips to the counterfactual answer (orange) in the final layers. This “flipping behavior” is most common when the model is prompted to respond with the world knowledge answer and provided with a counterfactual image (Figure 4 Panel B). This delayed integration of visual input leads to errors when the image contradicts memorized associations, matching the results in Table 1.
In contrast, when using “this” prompt (e.g., “what color is this strawberry?”) for identifying the counterfactual attribute, models are confident in the counterfactual answer by the middle layers and rarely flip to the world‐knowledge alternative (Figure 4 panel A and Table 2). This confidence is supported by the high inference accuracies seen in Table 1. Despite their confidence in the counterfactual answer, there is still a slight spike in world knowledge answer probability in mid-to-later layers. This slight spike shows that memorized knowledge does not fully disappear from the model, even when presented with contradicting inputs.
@@ 5.3 Controlling World Knowledge Associations with PvP Steering Vectors
LLaVA Color WK $\to$ CF 99.5 [14-16]
CF $\to$ WK 86.4 [10-17]
Size WK $\to$ CF 71.3 [8-16]
CF $\to$ WK 33.5 [12-21]
QwenVL Color WK $\to$ CF 99.7 [17-19]
CF $\to$ WK 78.8 [12-17]
Size WK $\to$ CF 89.9 [16-22]
CF $\to$ WK 61.8 [13-23]
Janus-Pro Color WK $\to$ CF 98.6 [14-16]
CF $\to$ WK 78.2 [15-18]
Size WK $\to$ CF 81.2 [16-19]
CF $\to$ WK 70.37 [16-20]
@@
Table 3 shows the effectiveness of our steering approach across tasks and models, highlighting both the percentage of successful steering of the model and the key layers at which intervention has the highest impact. We apply PvP steering vectors to the set of inputs that the model originally gets incorrect, meaning without PvP-steering, the model performance on this subset of data is 0%. Remarkably, we achieve at least a 98% success rate in flipping model predictions from world knowledge to counterfactual answers in the color task for all models. This demonstrates that MLLMs are not only steerable but highly responsive to targeted interventions, particularly when guided away from strongly encoded world knowledge priors. The size task is more difficult by nature: it requires detecting two objects and reasoning about their size relationship, making it more dependent on deeper, distributed visual processing. This is reflected in lower flip rates and broader intervention windows.
Across models, we observe that the most effective interventions tend to occur within specific mid-to-late layer ranges, typically requiring sustained influence over multiple layers (Table 3). In general, steering the model from world knowledge to a counterfactual (WK $\rightarrow$ CF) demands less intervention than reversing that shift (CF $\rightarrow$ WK), suggesting that overriding memorized priors is easier than restoring them once suppressed by a counterfactual input.
@@ 5.4 Impact of Prompting and Steering Vectors on Attention Mass
Figure 6: Impact of intervention on attention distribution. We observe more attention on text before intervention, and see a shift in attention towards the image and away from the text after intervention.
@@
Among all models, LLaVA-Next shows the strongest and most consistent shifts in attention when interventions are applied, followed by Qwen2-VL and then Janus-Pro. For the size task, steering vectors must act earlier in the network and across more layers to be effective, reflecting the fact that size requires integrating more visual features than color. In contrast, the color task is more localized and easier to influence with a smaller intervention window. Figure 5 illustrates how attention moves across layers in response to both prompt changes and steering task interventions in LLaVA-Next. To illustrate the same effect but with a concrete example, Figure 6, shows the attention before intervention being heavily focused on text, with “strawberries” highlighted in red (highest attention). After intervention, attention shifts to the image, with most red regions being inside the image, rather than over the text.
These findings show that PvP steering vectors reshape internal attention mechanisms more effectively than prompt changes alone. They offer precise control over how models allocate attention to visual inputs, especially in tasks like size comparison that require broader spatial reasoning. By intervening directly in the model’s representation space, PvP steering enables deeper interpretability and control over MLLM behavior.
@@ Appendix A Creating Visual CounterFact
(2) Size: Given two objects with the world knowledge relation “[object 1] is larger than [object 2]” we create the counterfactual image reflecting “[object 2] is larger than [object 1]” by combining the segmentation masks of object1 and object2. For the world knowledge image, we resize the masks so that object1 appears significantly larger than object2, and in the counterfactual image, we resize the masks so that object2 appears significantly larger than object1. Specifically, $\times250$ pixels versus $\times80$ pixels. This size difference visually reflects the intended relation.
@@ Appendix C Attention Mass Details
Color WK $\rightarrow$ CF $\% 12.812.8\%$
Color CF $\rightarrow$ WK $25.2\%$ $12.4\%$
@@
Color WK $\rightarrow$ CF $9.5\% 11$$1.7\%$
Color CF $\rightarrow$ WK $11.0\%$ $10.2\%$